# V-Zen: Efficient GUI Understanding and Precise Grounding With A Novel Multimodal LLM

### v1

@@ Abstract
In the rapidly evolving landscape of AI research and application, Multimodal Large Language Models (MLLMs) have emerged as a transformative force, adept at interpreting and integrating information from diverse modalities such as text, images, and Graphical User Interfaces (GUIs). Despite these advancements, the nuanced interaction and understanding of GUIs pose a significant challenge, limiting the potential of existing models to enhance automation levels. To bridge this gap, this paper presents V-Zen, an innovative Multimodal Large Language Model (MLLM) meticulously crafted to revolutionise the domain of GUI understanding and grounding. Equipped with dual-resolution image encoders, V-Zen establishes new benchmarks in efficient grounding and next-action prediction, thereby laying the groundwork for self-operating computer systems. Complementing V-Zen is the GUIDE dataset, an extensive collection of real-world GUI elements and task-based sequences, serving as a catalyst for specialised fine-tuning. The successful integration of V-Zen and GUIDE marks the dawn of a new era in multimodal AI research, opening the door to intelligent, autonomous computing experiences. This paper extends an invitation to the research community to join this exciting journey, shaping the future of GUI automation. In the spirit of open science, our code, data, and model will be made publicly available, paving the way for multimodal dialogue scenarios with intricate and precise interactions. Repo Link: SuperAGI/vzen
@@ 3 Proposed Architecture
Note: This is just a representational architecture. The final architecture diagram is yet to be drawn - It should contain the module’s names as per the equations in the paper
The architecture of V-Zen, our proposed multimodal Large Language Model (LLM), is a sophisticated ensemble of interconnected components meticulously designed for efficient GUI understanding and precise grounding. The architecture is composed of five major modules: Low-Resolution Visual Feature Extractor (LRVFE), Multimodal Projection Adapter (MPA), Pretrained Language Model with Visual Expert (PLMVE), High-Resolution Cross Visual Module (HRCVM), and High-Precision Grounding Module (HPGM).
@@ 3.4 High-Resolution Cross Visual Module
3.5 High-Precision Grounding Module
The HPGM takes the hidden states extracted from the PLMVE and uses them to perform precise grounding tasks [28, 35]. Unlike typical MLLM modules that provide grounding bounding boxes as part of the LLM’s text output, our HPGM outputs bounding box coordinates separately, ensuring precision. The module follows an enhanced DETR [3] object detector named DINO [17]. PLMVE’s last hidden state is used as the query of visual grounding to query the multi-scale feature set for visual grounding, denoted as $q_{llm\_gnd}$. The multi-scale feature set, denoted as $f{ms\_img}$, is obtained using a Swin Transformer-based backbone. It takes $q_{llm\_gnd}$ and $f{ms\_img}$ and produces the bounding boxes for precise grounding. This way, the HPGM module can precisely ground the GUI elements based on the processed image and text features.
@@ 3.5 High-Precision Grounding Module
4 Experiments and Results: To Be Re-Written Later

### v2

@@ Abstract
In the rapidly evolving landscape of AI research and application, Multimodal Large Language Models (MLLMs) have emerged as a transformative force, adept at interpreting and integrating information from diverse modalities such as text, images, and Graphical User Interfaces (GUIs). Despite these advancements, the nuanced interaction and understanding of GUIs pose a significant challenge, limiting the potential of existing models to enhance automation levels. To bridge this gap, this paper presents V-Zen, an innovative Multimodal Large Language Model (MLLM) meticulously crafted to revolutionise the domain of GUI understanding and grounding. Equipped with dual-resolution image encoders, V-Zen establishes new benchmarks in efficient grounding and next-action prediction, thereby laying the groundwork for self-operating computer systems. Complementing V-Zen is the GUIDE dataset, an extensive collection of real-world GUI elements and task-based sequences, serving as a catalyst for specialised fine-tuning. The successful integration of V-Zen and GUIDE marks the dawn of a new era in multimodal AI research, opening the door to intelligent, autonomous computing experiences. This paper extends an invitation to the research community to join this exciting journey, shaping the future of GUI automation. In the spirit of open science, our code, data, and model will be made publicly available, paving the way for multimodal dialogue scenarios with intricate and precise interactions. Repo Link: github.com/abdur75648/V-Zen
@@ 3 Proposed Architecture
The architecture of V-Zen, our proposed multimodal Large Language Model (LLM), is a sophisticated ensemble of interconnected components meticulously designed for efficient GUI understanding and precise grounding. The architecture is composed of five major modules: Low-Resolution Visual Feature Extractor (LRVFE), Multimodal Projection Adapter (MPA), Pretrained Language Model with Visual Expert (PLMVE), High-Resolution Cross Visual Module (HRCVM), and High-Precision Visual Grounding Module (HPVGM).
@@ 3.4 High-Resolution Cross Visual Module
3.5 High-Precision Visual Grounding Module
The HPVGM takes the hidden states extracted from the PLMVE and uses them to perform precise grounding tasks [28, 35]. Unlike typical MLLM modules that provide grounding bounding boxes as part of the LLM’s text output, our HPVGM outputs bounding box coordinates separately, ensuring precision. The module follows an enhanced DETR [3] object detector named DINO [17]. PLMVE’s last hidden state is used as the query of visual grounding to query the multi-scale feature set for visual grounding, denoted as $q_{llm\_gnd}$. The multi-scale feature set, denoted as $f{ms\_img}$, is obtained using a Swin Transformer-based backbone. It takes $q_{llm\_gnd}$ and $f{ms\_img}$ and produces the bounding boxes for precise grounding. This way, the HPVGM module can precisely ground the GUI elements based on the processed image and text features.
@@ 3.5 High-Precision Visual Grounding Module
4 Experiments and Results